### v2

# Probing Non-Fermi-Liquid Behaviour of Composite Fermi Liquid via Efficient Thermal Simulations

Abstract
The physics of two-dimensional electron gas in a perpendicular magnetic field, i.e., the quantum Hall system, is remarkably rich. At half filling of the lowest Landau level, it has been predicted that “composite fermions”—emergent quasiparticle of an electron with two magnetic flux quanta—experience zero net magnetic field and form a Fermi sea, dubbed composite Fermi liquid (CFL). However, the seemingly simple appearance of CFL is a strongly correlated quantum many-body state in disguise, and to solve it in a controlled manner is extremely difficult, to the level that the thermodynamic properties of CFL is still largely unknown. In this work, we perform state-of-the-art thermal tensor network simulations on the $\nu=1/2$ Landau level systems, and observe low-temperature power-law behaviour of the specific heat, signaling the gapless nature of CFL. More importantly, the power is extracted to be closed to $2/3$, clearly deviated from the ordinary linear-$T$ Fermi liquid behaviour, suggesting the coupling between the CFs and the dynamical emergent gauge field and therefore revealed the quantum many-body aspect of the CFL state. Relevance of our methodology to other quantum Hall settings and moiré systems is discussed.
Introduction.— The two-dimensional electron gas subjected to a perpendicular magnetic field—known as the quantum Hall (QH) system—exhibits remarkable physical richness. In the integer quantum Hall regime, the Hall conductance $\sigma_{xy}$ is quantized precisely in integer multiples of $e^{2}/h$ [1]. In the fractional quantum Hall regime, $\sigma_{xy}$ quantizes at fractional values of $e^{2}/h$, accompanied by the emergence of fractionally charged quasiparticles obeying abelian fractional statistics [2, 3, 4]. Notably, at filling $\nu=1/2$, the system exhibits metallic behavior, characterized by a finite longitudinal resistivity $\rho_{xx}$ and the absence of $\rho_{xy}$ pleateau  [5].
The composite fermion (CF) theory postulates that electrons attach an even number of magnetic flux quanta, forming emergent quasiparticles that experience a reduced effective magnetic field [6, 7]. Along the line of this flux-attachment concept, mean-field theories [8, 9, 10, 11] are successful: near $\nu=1/2$, ballistic transport experiments are consistent with composite fermions moving in a weak residual magnetic field [12, 13, 14]; at half-filling, the external magnetic field is entirely canceled by the attached flux, resulting in a Fermi sea of composite fermions—a picture corroborated by surface acoustic wave measurements [15]. More recently, this conventional CF picture has been revisited in light of connections to 3D topological insulators (TIs), reformulating composite fermions as massless Dirac fermions reminiscent of TI surface states [16, 17, 18, 19]. This “composite Dirac fermion” theory couples fermions to an emergent gauge field, exhibiting non-Fermi-liquid behavior and providing a field-theoretic realization of (2+1)-dimensional quantum electrodynamics (QED3) as a strongly coupled conformal field theory, with both condensed matter [20, 21, 22, 23, 24, 25, 26, 27, 28] and high-energy [29, 30, 31, 32, 33, 34, 35, 36] physics correspondence.
In recent years, the fractional Chern insulator (FCI) [37, 38, 39, 40, 41, 42]—a lattice analogue of the fractional quantum Hall effect (when there is no mangetic field the state is also denoted as fractional quantum anomalous Hall (FQAH) effect)—has attracted considerable interest in the context of two-dimensional moiré materials, with extensive theoretical explorations [43, 44, 45, 46, 47, 48, 49, 50, 51]. FCIs are reported at high magnetic field in early experiment [52] at the fractional filling of Hofstadter band [53], and the fractional filling of the native Chern band in magic-angle twisted bilayer graphene (TBG) with the finite magnetic field to improve the quantum geometry, i.e., smoothing the Berry-curvature distribution and reducing violations of the trace condition of the ideal Landau-level-like relation $\operatorname{Tr}(g(\mathbf{k}))=|\Omega(\mathbf{k})|$ [54, 55]. More recently, the zero-field FCIs, i.e. FQAHs, are observed in twisted bilayer MoTe2 (tMoTe2) and rhombohedral graphene/hBN superlattices [56, 57, 58, 59, 60]. Notably, the zero-field metallic state at $\nu=1/2$ of the topological bands have also been theoretically and experimentally explored [61, 62, 57, 60].
Substantial numerical research are dedicated to understanding the gapped FQH as well as the zero-field FCI phases, including the time-dependent variational principle [63, 64, 65] and single mode approximation [66] concerning the spectral properties, and thermal tensor network simulations [67] on the thermodynamical properties. In contrast, the numerical characterization of the half-filled CFL remains challenging. From this perspective, representative works include density matrix renormalization group (DMRG) investigation on the role of particle-hole symmetry  [68], variational Monte Carlo studies on the scaling of entanglement entropy [69, 70, 71], effects of anisotropy [72], DMRG studies on the relationship between the Fermi sea of bare electrons at zero field and that of composite fermions at high field [73] as well as the influence of discrete rotational symmetry $C_{N}$ [74], and exact diagonalization calculation on the thermoelectric response [75]. In all these directions, new computation and discussions of often-time contradicting results are actively going on.
In this work, we investigate the thermodynamic properties of the composite Fermi liquid at half filling of the lowest Landau level using large-scale thermal tensor network simulations, specifically the tangent-space tensor renormalization group (tanTRG) method [76]. We observe a low-temperature specific heat that scales as $T^{\alpha}$, with $\alpha\approx 2/3$—indicative of gapless excitations and markedly inconsistent with the linear-$T$ dependence expected for a conventional Fermi liquid. This power-law behavior provides direct evidence for the strongly correlation nature of CFL state, in that, the coupling between composite fermions and a dynamical emergent gauge field gives rise to non-Fermi liquid properties. In a broader context, our results put forward the theoretical framework that such a non-Fermi-liquid could emerge from Fermi surface coupled to a critical ($U(1)$) gauge field [25, 26], ferromagnetic bosonic fluctuations [77, 78] and nematic fluctuations [79]. And our thermal tensor-network method can be straightforwardly generalized to FQAH and moiré systems, where the continuum Hamiltonian is naturally defined in momentum space.
Figure 1: The matrix product operator (MPO) setup in the lowest Landau level (LLL) basis for thermal density operator $\rho$. (a) In the Landau gauge with periodic boundary condition along the $y$-direction and open boundary condition along the $x$-direction. The LLL wavefunctions (labelled by $n\in[0,N-1]$) are extended along $y$- and localized along $x$-direction. (b) The matrix product operator of density operator $\rho$ is defined in the LLL basis, where the “local” degree of freedom of the $i$th tensor lives in the 2-dimensional physical space of either filling the $i$th LLL or leaving it empty.
Model definition.— The model Hamiltonian we considered is defined as
$\mathcal{H}\equiv\int\mathrm{d}\mathbf{r}_{1}\mathrm{d}\mathbf{r}_{2}~V(\mathbf{r}_{1},\mathbf{r}_{2})[n(\mathbf{r}_{1})-\tfrac{1}{2}][n(\mathbf{r}_{2})-\tfrac{1}{2}],$ (1)
where $V(\mathbf{r}_{1},\mathbf{r}_{2})\equiv\frac{e^{-|\mathbf{r}_{1}-\mathbf{r}_{2}|/\lambda}}{|\mathbf{r}_{1}-\mathbf{r}_{2}|}$ is the Coulomb-Yukawa interaction between electrons, $n(\mathbf{r})\equiv c^{\dagger}(\mathbf{r})c(\mathbf{r})$ is the particle number operator and the factor $\tfrac{1}{2}$ is utilized to keep the half-filling condition in the grand-canonical ensemble. We then project the Hamiltonian onto the lowest Landau level (LLL) via $c(\mathbf{r})=\sum_{n}\phi_{n}(\mathbf{r})c_{n}$, where $\phi_{n}(\mathbf{r})\propto e^{-\frac{1}{2}(\frac{x-x_{n}}{\ell})^{2}}e^{\mathrm{i}k_{n}y}$ is the LLL wavefunction in the cylindrical geometry (i.e., Landau gauge) with $k_{n}=2\pi n/L_{y}$, $x_{n}=2\pi n\ell^{2}/L_{y}$, $n\in[0,N-1]$ and $N$ the number of LLL orbitals [c.f. Fig. 1 (a)]. We set the magnetic length $\ell=\sqrt{\hbar c/(eB)}=1$ and $k_{\mathrm{B}}=1$ as the unit throughout. This yields the LLL-projected Hamiltonian,
$\mathcal{H}^{\mathrm{LLL}}=\sum_{m,n,k,l}\mathcal{A}_{m,n,k,l}c^{\dagger}_{m}c^{\dagger}_{n}c^{\,}_{k}c^{\,}_{l},$ (2)
with the form factor $\mathcal{A}_{m,n,k,l}=\int\mathrm{d}r_{1}\mathrm{d}r_{2}V(r_{1}-r_{2})\phi^{\ast}_{m}(r_{1})\phi_{l}(r_{1})\phi^{\ast}_{n}(r_{2})\phi_{k}(r_{2})$ being evaluated numerically.
Tangent-space tensor renormalization group (tanTRG) in the LLL basis.— Here, we briefly recapitulate the idea of tangent-space tensor renormalization group (tanTRG) which is proved to be successful in accessing the challenging low-temperature properties of quantum lattice models, e.g., 2D Hubbard model and $t$-$J$ model for cuprate [76, 80], bilayer $t$-$J$-$J_{\perp}$ model for Nickelate superconductor [81], and real-space effective model of MoTe2 [82]. We then show how tanTRG can be adapted to the LLL orbital basis.
To begin with, a generic Hamiltonian can be efficiently decomposed into matrix product operator (MPO) representation [83, 84, 85, 86, 87]. At high temperature, the themal density operator can then be constructed via series expansion with high precision [88], which effectively reduce to $\rho(\tau)\equiv e^{-\tau\mathcal{H}}=I-\tau\mathcal{H}$ for sufficiently small inverse temperature $\tau\equiv 1/T$ (here $\tau=10^{-6}$ is taken rendering a small expansion error of $\mathcal{O}(10^{-12})$).
We then cool down the system through a mixed temperature grid, i.e., firstly a exponential one $\beta_{i}=\tau\times 2^{i/z}$ to quickly approaching the low-$T$ regime (here $z=4$ to ensure dense temperature point for numerical derivative), which then follows by a linear (inverse) temperature grid when the inverse temperature increase $\Delta\beta$ exceeds a given threshold (here $\Delta\beta=5$ is taken) to ensure small projection errors in low-$T$ regime.
The cooling step is followed by the flow equation $\partial\rho/\partial\beta=\mathcal{P}(-\mathcal{H}\rho)\mathcal{P}$ constraint to a fixed-bond-dimension matrix product operator (MPO) manifold $\mathcal{M}_{\mathrm{MPO}}$, with $\mathcal{H}\rho$ being the tangent vector in the full Hilbert space which generally takes the density matrix $\rho$ out of $\mathcal{M}_{\mathrm{MPO}}$, and the projector $\mathcal{P}$ serves to projecting the tangent vector back to the tangent space of the MPO manifold $\mathcal{T}_{\rho}\mathcal{M}_{\mathrm{MPO}}$. For more details concerning the MPO parameterization of the above constraint flow equation, we refer the interested readers to Ref. [76].
Specifically, in this work, we deal with the many-body basis consisted of LLL wavefunctions, $|q_{0},q_{1},\cdots,q_{N-1}\rangle=(c^{\dagger}_{0})^{q_{0}}(c^{\dagger}_{1})^{q_{1}}\cdots(c^{\dagger}_{N-1})^{q_{N-1}}|\Omega\rangle$, with the quantum number $q_{i}$ denoting the number of particles filled in the $i$th LLL orbital, and $|\Omega\rangle$ being the vacuum state. A generic operator $O$ in this many-body LLL basis then express as $O=\sum_{\{q_{i},q_{i}^{\prime}\}}O_{\{q_{i}\},\{q_{i}^{\prime}\}}|\{q_{i}\}\rangle\langle\{q_{i}^{\prime}\}|$, where $O_{\{q_{i}\},\{q_{i}^{\prime}\}}\equiv\langle\{q_{i}\}|O|\{q_{i}^{\prime}\}\rangle$ is a $2^{N}\times 2^{N}$-dimensional tensor, and can be parameterized into the MPO form, $O_{\{q_{i}\},\{q_{i}^{\prime}\}}=\sum_{b_{0},b_{1},\cdots,b_{N}}\prod_{i=0}^{N-1}A_{b_{i},b_{i+1},q_{i},q_{i}^{\prime}}$ with the virtual indices $b_{i}$’s ranging from $1$ to $D$ (bond dimension), as shown in Fig. 1 (b). For the projected Hamiltonian $\mathcal{H}^{\mathrm{LLL}}$ and the initial high-temperature density operator $\rho(\tau)$, we can find the exact MPO forms for them, and thus the lower-temperature density operator $\rho(\beta)$ can then be tracked with the tanTRG protocol as mentioned above. We note that our method can be straightforwardly generalized to FQAH and quantum morié systems [89, 43, 44, 45, 46, 47, 48, 49, 50, 51, 90] and other projected Hamiltonian [91, 92] systems, where one only needs to change the basis function on which the projection is carried out.
Figure 2: Benchmark and CBE optimization of tanTRG for the projected Coulomb-Yukawa interaction. We chose $N=16$ and compared to the exact diagonalization results. (a) Relative errors of free energy $\delta f\equiv|f_{\mathrm{tanTRG}}-f_{\mathrm{ED}}|/|f_{\mathrm{ED}}|$ calculated with different bond dimensions $D=400,800,1200$ are shown versus temperature $T$. (b) Relative errors of specific heat $\delta c_{\mathrm{V}}$ are shown versus temperature $T$. In the inset of both panels, the raw tanTRG data of the free energy $f$ and specific heat $c_{\mathrm{V}}$ is shown along with the ED data (the yellow curves). (c) Relative errors of specific heat $\delta c_{\mathrm{V}}$ for both 2-site (blue, the same as in (a)) and CBE (red) update scheme with various bond dimensions $D=400,800,1600$. (d) The CPU time (in the unit of hours) for both 2-site and CBE update scheme, shown as functions of bond dimension $D$.
Benchmark results.— In Fig. 2, we benchmark tanTRG in a small system size $N=16$. Here, we take the circumference of the cylinder $L_{y}=12$, and the effective screening length $\lambda=10\tfrac{2\pi\ell^{2}}{L_{y}}$. We firstly calculate the free energy per orbital $f(T)\equiv-\frac{1}{N}T\ln Z$ with $Z=\operatorname{Tr}\rho$ the partition function.
In Fig. 2 (a), we have shown three curves of the free energy relative errors $\delta f\equiv|f_{\mathrm{tanTRG}}-f_{\mathrm{ED}}|/|f_{\mathrm{ED}}|$ with incremental bond dimensions $D=400,800,1600$. For all the three curves, at the high-temperature regime, the relative errors remain within the order of $10^{-6}$ as the discarded weights can be neglected and the final error comes from the Trotter error and the projection error (the one results from projecting the full tangent vector into $T_{\rho}\mathcal{M}_{\mathrm{MPO}}$). At intermediate temperature $T\sim 10^{-1}$, the thermal state changes significantly as indicated by the round peak of the specific heat [c.f. the inset of Fig. 2 (b)], and the finite discarded weights set in, resulting in a relatively higher errors of free energies. At lower-temperature regime, finite-size gap sets in and the thermal density operator $\rho$ converges to $|\psi_{0}\rangle\langle\psi_{0}|$ (ground state) which in this example possess lower (purified) entanglement entropy thus lower $\delta f$ relative to the intermediate-$T$ cases. Such lower-$T$ behaviour can also be understood from the MPO manifold perspective, which becomes more flattened when the finite-size gap sets in and renders small projection errors. Furthermore, by increasing $D$, the precision increases from $10^{-3}$ to $10^{-5}$ accordingly in the low-$T$ regime.
In Fig. 2 (b), we present the relative error of specific heat $c_{V}\equiv\partial e/\partial T$, with $e\equiv\frac{1}{N}\operatorname{Tr}(\mathcal{H}\rho)/Z$ being the internal energy per orbital. Compared with the exact diagonalization results, the relative errors of specific heat $\delta c_{V}$ is shown as function of temperature $T$. Similar to the free energy, the specific heat $c_{V}$ possesses high precision with relative errors within $10^{-8}$ for high temperatures $T\gtrsim 1$, which increase quickly at $T\sim 10^{-1}$, and at lower temperatures $\delta c_{V}$ even decreases slightly due to existence of the finite-size gap. Furthermore, by increasing bond dimensions $D=400\to 1600$, the precision again gets improved from $10^{-2}$ to $10^{-4}$.
The computational complexity can be further reduced by replacing the 2-site (2S) update scheme with the recently-proposed controlled bond expansion (CBE) [93, 94], which has shown improved computational efficiency while keeping similar precision as the 2S scheme in DMRG and TDVP simulations. Theoretically, this will result in a speed up by a factor of $d$ (here $d=2$) in MPS based algorithms, and by a factor of $d^{2}$ in MPO based algorithms which is very promising for our finite-$T$ simulations.
In Fig. 2 (c), we show the relative errors of the specific heat with respect to the ED results. It can be seen that, at low-temperature regime $T\lesssim 10^{-1}$, the 2S and CBE results exhibit similar accuracies for various bond dimensions from 400, 800, to 1600, which proves its validity in the considered projected Hamiltonian. In Fig. 2 (d), we compare the computational complexity of these two types of simulations (2S and CBE). Overall, the CBE simulations consume much less CPU time compared to the 2S ones. At small bond dimensions, due to the numerical overheads, the acceleration of CBE scheme is below the ideal factor $d^{2}=4$. But as the bond dimensions increase, the speed up effect of CBE is getting increasingly pronounced; i.e., at the largest bond dimension $D=1600$ consindered, the acceleration is over a factor of $3$. Overall, we find tanTRG combined with CBE update provides an efficient and high-precision approach to the study of the finite-temperature properties of the LLL-projected Hamiltonians.
Figure 3: Thermodynamic properties of composite Fermi liquid. (a) Lower-temperature specific heat $c_{\mathrm{V}}$ behaviour for various number of orbitals $N=24,36,48$, exhibiting converged power-law scaling $T^{2/3}$ for $N=36$ and 4848. (b) The low-$T$ power $\alpha_{\mathrm{Fit}}$ from a linear fitting is shown versus system sizes $1/N$ for $\lambda=10$ and 22.
Figure 4: Formation of the composite Fermi surface in CFL. For the system with $L_{y}=12,N=48,\lambda=10$, (a) Two-dimensional momentum space view of the “guiding center” density-density correlations $\bar{D}(q)=\exp(q^{2}/2)\langle n_{q}n_{-q}\rangle_{T}$ is plotted as a function of 2D momenta $q=(q_{x},q_{y})$ at four different temperatures $T=10^{1},10^{0},10^{-1},10^{-2}$. It starts to exhibit a $2k_{F}$ circle for $T\lesssim 10^{-1}$ suggesting the establishment of the composite Fermi surface. (b,c) $\bar{D}(q)$ is shown as functions of $q_{x}$ when $q_{y}=0,\frac{2\pi}{L_{y}}$ are fixed accordingly. The dashed grey line indicated the $2k_{F}$ points for the corresponding $q_{y}$. (d) $\bar{D}(q)$ at the three different $2k_{F}$ points of $q_{y}=m\frac{2\pi}{L_{y}}$ with $m=0,1,2$, which increase most quickly at $\sim 10^{-1}$, suggesting the formation of the composite Fermi surface.
Finite-temperature properties of Composite Fermi Liquid.— We then move forward to larger system sizes with various number of orbitals $N=24,36,48$ with a fixed circumference $L_{y}=12$ and a relatively long screening length $\lambda=10$. As shown in Fig. 3 (a), we calculate the specific heat $c_{V}$ down to low temperatures $T\lesssim 10^{-2}$. For the following analysis, we exclude the lower-$T$ regime data as depicted by the grey area to minimize the finite-size effect. This is suggested by the compressibility $\partial\langle n\rangle/\partial\mu$ data [c.f. SM [95]] which quickly vanishes at $\sim 10^{-2}$, indicating a finite-circumference gap. Above the grey regime, the three specific heat curves exhibit power-law behaviour $T^{\alpha}$, and they all apparently deviate from the linear-$T$ Fermi-liquid behaviour. As shown in Fig. 3 (b), notably, for the two larger sizes $N=36$ and $48$, the power $\alpha$ converge to the value of $2/3$, consistent with the value predicted by Halperin, Lee, and Read (HLR) [11]. In the plot, we have also shown the smaller $\lambda=2$ data, they seem to suffer severer finite-size effect, while the power still tend to approach to $2/3$. The fact that such a non-Fermi-liquid thermodynamic response – different from the $C\sim T$ of normal metal, is the decisive evidence that there exists strong interactions among the CFs, and such interactions are mediated by the emergent gauge field in the QED3 description [11, 16, 17, 18, 19, 25, 26], or ferromagnetic and nematic critical bosonsic fluctuations [79, 77, 78].
We can also demonstrate the formation of the CFL state during the cooling process. In the 2D limit, the composite Fermion in CFL are expected to form a circular Fermi surface with Fermi momentum $k_{F}=1$. But it is not manifested in the spectral function of the bare electrons, rather, due to the non-Fermi-liquid nature of the state, can be verified by the “guiding center” density-density correlation $\bar{D}(q)=\exp(q^{2}/2)\langle n_{q}n_{-q}\rangle_{T}=\exp(q^{2}/2)\operatorname{Tr}[n_{q}n_{-q}\exp(-\mathcal{H}/T)]/Z$, which reflects the scattering events of CFs near the composite Fermi surface with a $2k_{F}$ singularity circle in the momentum space. In Fig. 4 (a), we plot $\bar{D}(q)$ as a function of 2D momenta $q=(q_{x},q_{y})$ at four different temperatures $T=10^{1},10^{0},10^{-1},10^{-2}$. It starts to exhibit a $2k_{F}$ circle for $T\lesssim 10^{-1}$, i.e. the evidence for the formation of the composite Fermi surface. As shown in the panel (b) and (c), we choose two fixed-$q_{y}$ (specifically, $q_{y}=0,2\pi/L_{y}$) cuts in the 2D momentum contour. The dashed grey lines indicate the $2k_{F}$ points for the corresponding $q_{y}$’s. One can then see a clear round peak around the $2k_{F}$ points, which will become the true singularity points in the 2D limit, and the round peaks build up significantly around $T\sim 10^{-1}$ manifesting our finding above. This can also be seen in the panel (d), where $\bar{D}(q)$ at the three different $2k_{F}$ points of $q_{y}=m\frac{2\pi}{L_{y}}$ with $m=0,1,2$, which increase most quickly at $\sim 10^{-1}$. Such temperature scale matches with that of non-Fermi-liquid specific heat observed in Fig. 3. We note that, in the low-temperature regime, the small-$q$ density correlation $D(q)$ shows a Fermi-liquid $q^{3}$ behaviour [c.f. SM [95]]. This implies the CFs by themselves, to some extent, behave as electrons in conventional Fermi liquid, and it is the strong coupling with the emergent gauge field that contributes to the non-Fermi-liquid behaviour of the specific heat.
[…]
Furthermore, our findings demonstrate the efficacy of the thermal tensor network approach within the framework of projected Hamiltonian [91, 92], particularly in momentum space, highlighting its potential utility in continuum models in the context of other FQH systems [96] and FQAH and quantum moiré materials – only need to change the basis function – to study the thermodynamics, magnetism and topology [97, 98, 99]. E.g., our methodological can be straightforwardly applied to the zero-field CFL in the twisted TMD system [61, 62, 57, 60] to explore whether such a sublinear-$T$ specific heat still holds in the absence of Landau level, which might offer a thermal signature that distinguish the CFL from Fermi liquid in the quantum moiré experiments.
Acknowledgment.- We thank Dung Xuan Nguyen, Bo Yang and Duncan Haldane for discussion on the subject. We acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17309822, C7037-22GF, 17302223, 17301924, 17301725), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank HPC2021 system under the Information Technology Services at the University of Hong Kong [100], as well as the Beijing Paratera Tech Corp., Ltd [101] for providing HPC resources that have contributed to the research results reported within this paper.
SUPPLEMENTARY MATERIALS FOR
PROBING NON-FERMI-LIQUID BEHAVIOUR OF COMPOSITE FERMI LIQUID VIA EFFICIENT THERMAL SIMULATIONS
In Supplementary Materials Section I, we provide detailed derivation of a generic interaction (e.g. Coulomb-Yukawa interaction considered in the main text) projected onto the lowest Landau level on a cylinder. In Section II, we provided more detailed thermodynamic data of composite Fermi Liquid at the half filling of lowest Landau level.
Section I Generic Interaction projected onto LLL basis
We consider a generic two-body interaction $V(r_{1}-r_{2})$,
$\mathcal{H}$ $=\int dr_{1}\int dr_{2}V(r_{1}-r_{2}):\hat{\rho}(r_{1})\hat{\rho}(r_{2}):$ (S1)
$=\int dr_{1}\int dr_{2}V(r_{1}-r_{2}):c^{\dagger}(r_{1})c(r_{1})c^{\dagger}(r_{2})c(r_{2}):$ (S2)
where $c(r)$ is a spinless fermionic annihilation operator on position $r$. We then project it onto the lowest Landau level on a $L_{x}\times L_{y}$ cylinder (periodic boundary condition along $y$ axis), i.e.,
$c(r)=\sum_{n}\phi_{n}(r)c_{n}$ (S3)
Here, we take the Landau gauge $(A_{x},A_{y})=(0,Bx)$, and the wavefunction takes the form of
$\phi_{n}(r)=\frac{1}{\pi^{1/4}\sqrt{L_{y}\ell}}e^{-\frac{1}{2}(\frac{x-x_{n}}{\ell})^{2}}e^{ik_{n}y},$ (S4)
with $k_{n}=\frac{2\pi n}{L_{y}}$, $x_{n}=\frac{2\pi n}{L_{y}}\ell^{2}$, and the magnetic length $\ell=\sqrt{\hbar c/(eB)}$. We then arrive at
$\mathcal{H}$ $=\sum_{n_{1},m_{1},n_{2},m_{2}}\int dr_{1}\int dr_{2}V(r_{1}-r_{2})\phi^{\ast}_{n_{1}}(r_{1})\phi_{m_{1}}(r_{1})\phi^{\ast}_{n_{2}}(r_{2})\phi_{m_{2}}(r_{2}):c^{\dagger}_{n_{1}}c_{m_{1}}c^{\dagger}_{n_{2}}c_{m_{2}}:$ (S5)
$=\sum_{n_{1},m_{1},n_{2},m_{2}}\mathcal{A}_{n_{1},n_{2},m_{2},m_{1}}:c^{\dagger}_{n_{1}}c_{m_{1}}c^{\dagger}_{n_{2}}c_{m_{2}}:$ (S6)
For the form factor, we have
$A_{n_{1},n_{2},m_{2},m_{1}}=\int dr_{1}\int dr_{2}V(r_{1}-r_{2})\phi^{\ast}_{n_{1}}(r_{1})\phi_{m_{1}}(r_{1})\phi^{\ast}_{n_{2}}(r_{2})\phi_{m_{2}}(r_{2}).$ (S8)
For the Yukawa interaction
$V(r)=\frac{e^{2}}{4\pi\epsilon}\frac{1}{r}\exp(-r/\lambda),$ (S9)
with the Fourier transformation
$\tilde{V}(Q)=\frac{e^{2}}{4\pi\epsilon}\frac{1}{\sqrt{Q^{2}+1/\lambda^{2}}}=\frac{e^{2}}{4\pi\epsilon}\frac{1}{\sqrt{q_{x}^{2}+q_{y}^{2}+1/\lambda^{2}}}.$ (S10)
The form factor matrix element will then be
$\mathcal{A}_{n_{1},n_{2},n_{3},n_{4}}=$ $2\frac{e^{2}}{4\pi\epsilon}\frac{\delta_{n_{1}+n_{2},n_{3}+n_{4}}}{L_{y}}\exp\left(-\frac{1}{2}(k_{1}-k_{4})^{2}\right)\int_{0}^{\infty}dq_{x}\frac{\cos(q_{x}(k_{1}-k_{3}))}{\sqrt{q_{x}^{2}+(k_{1}-k_{4})^{2}+1/\lambda^{2}}}\exp\left(-\frac{1}{2}q_{x}^{2}\right).$ (S11)
Section II More data on the thermodynamics of CFL
Figure S1: Charge compressibility $\partial\langle n\rangle/\partial\mu$ is shown versus $T$, which quickly drops to zero at the low-$T$ regime (denoted by the grey area) suggesting the finite-size effect.
Figure S2: Guiding-center density correlation $\bar{D}(q)$ data at low-$T$ regime. (a) For $\lambda=10$ and $N=48$, $\bar{D}(q)$ is shown versus $q_{x}$ at fixed $q_{y}=0$, showing power-law behaviour $q_{x}^{\alpha}$ in low-$T$ and small-$q$ regime. At relatively low temperature $T=0.0055$, the extracted $\alpha$ is $3.01(6)$, close to the Fermi-liquid behaviour $q^{3}$. (b) The extracted $\alpha_{\mathrm{Fit}}$ is shown versus temperature $T$, which converges to $3$ at low-$T$ regime within the errorbar of the linear fitting. (c,d) Similar plots as (a,b) with smaller $\lambda=2$, where the low-$T$ behaviour is slightly deviated from Fermi-liquid behaviour $q^{3}$.
As shown in Fig. S1, for fixed circumference $L_{y}=12$ and $\lambda=10$, the charge compressibility $\partial\langle n\rangle/\partial\mu$ is calculated by including a chemical potential term $\mu\sum_{n}\hat{\rho}_{n}$ in the model Hamiltonian Eq. (2), and thus via a grand-canonical-ensemble simulations. The obtained charge compressibility is shown to quickly vanish at $T\sim 10^{-2}$ signaling that, the system enters an incompressible state due to the finite-size effect. Below $T\sim 10^{-2}$, the thermodynamics data, e.g. the specific heat data, is thus excluded from the analysis in the main text.
As shown in Fig. S2, we have calculated the guiding-center density correlation $\bar{D}(q)$ data at low-$T$ regime. In panel (a), for $\lambda=10$ and $N=48$, $\bar{D}(q)$ is shown versus $q_{x}$ at fixed $q_{y}=0$, showing power-law behaviour $q_{x}^{\alpha}$ in low-$T$ and small-$q$ regime. At relatively low temperature $T=0.0055$, the extracted $\alpha$ is $3.01(6)$, close to the Fermi-liquid behaviour $q^{3}$. In panel (b), the extracted $\alpha_{\mathrm{Fit}}$ is shown versus temperature $T$, which converges to $3$ at low-$T$ regime within the errorbar of the linear fitting. As shown in panels (c) and (d), for a smaller $\lambda=2$, similar to the specific heat behaviour, it seems to suffer severe finite-size effect as the extracted power $\alpha$ oscillates around $3$ at low-$T$ regime.